need help please.

Examine the balance sheet of commercial banks in the following table.

Assets $ Billion % Total Liabilities and Net Worth $ Billion % Total
Real assets Liabilities
Equipment and premises $

193.7

1.3 % Deposits $ 10,277.3 67.1 %
Other real estate 48.9 0.3 Debt and other borrowed funds

990.0

6.5
Total real assets $ 242.6 1.6 % Federal funds and repurchase agreements 942.0 6.2
Other 929.0 6.1
Total liabilities $ 13,138.3 85.8 %
Financial assets
Cash $ 890.3 5.8 %
Investment securities

5,032.1

32.9

Loans and leases

6,711.3

43.8
Other financial assets

1,215.2

7.9

Total financial assets $ 13,848.9 90.5 %
Other assets
Intangible assets $

423.4

2.8 %
Other 794.7 5.2
Total other assets $

1,218.1

8.0

% Net worth $ 2,171.3 14.2 %
Total $ 15,309.6 100.0 % $ 15,309.6 100.0 %

Balance sheet of FDIC-insured commercial banks and savings institutions

Note: Column sums may differ from total because of rounding error.

Source: Federal Deposit Insurance Corporation, www.fdic.gov, March 2016.

a. What is the ratio of real assets to total assets? (Round your answer to 4 decimal places.)

b. What is the ratio of real assets to total assets for nonfinancial firms in the following table? (Round your answer to 4 decimal places.)
Answer #1

A) The ratio of real assets to total assets is calculated as

242.6/15309.6 =.0158

b) ratio of real assets to other assets of non financial firm = 242.6/15067=.01610